# Agentless [Uncaptioned image] : Demystifying LLM-based Software Engineering Agents

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly advanced the automation of software development tasks, including code synthesis, program repair, and test generation. More recently, researchers and industry practitioners have developed various autonomous LLM agents to perform end-to-end software development tasks. These agents are equipped with the ability to use tools, run commands, observe feedback from the environment, and plan for future actions. However, the complexity of these agent-based approaches, together with the limited abilities of current LLMs, raises the following question: Do we really have to employ complex autonomous software agents? To attempt to answer this question, we build Agentless – an agentless approach to automatically solve software development problems. Compared to the verbose and complex setup of agent-based approaches, Agentless employs a simplistic two-phase process of localization followed by repair, without letting the LLM decide future actions or operate with complex tools. Our results on the popular SWE-bench Lite benchmark show that surprisingly the simplistic Agentless is able to achieve both the highest performance (27.33%) and lowest cost (\$0.34) compared with all existing open-source software agents! Furthermore, we manually classified the problems in SWE-bench Lite and found problems with exact ground truth patch or insufficient/misleading issue descriptions. As such, we construct SWE-bench Lite-$S$ by excluding such problematic issues to perform more rigorous evaluation and comparison. Our work highlights the current overlooked potential of a simple, interpretable technique in autonomous software development. We hope Agentless will help reset the baseline, starting point, and horizon for autonomous software agents, and inspire future work along this crucial direction. We have open-sourced Agentless at: https://github.com/OpenAutoCoder/Agentless
@@ 1 Introduction
Large language models (LLMs) have become the go-to default choice for code generation [Chen et al., 2021; Austin et al., 2021; Li et al., 2023; Wei et al., 2023]. State-of-the-art LLMs like GPT-4 [OpenAI, 2023] and Claude-3.5 Anthropic [2024] have demonstrated their prowess in being able to synthesize code snippets based on given user description. However, compared to the main evaluation setting of simple, self-contained problems, applying LLMs on repository-level software engineering tasks has been understudied. Software engineering tasks like feature addition, program repair, and test generation require an in-depth understanding of not only information within files, which can contain thousands of lines of code, but also repository-level dependencies across files.
@@
To solve the challenging real-world software development problems from SWE-bench, inspired by the Devin AI Software Engineer dev [2024], there has been a significant body of work from both academia and industry focusing on developing agent-based approaches Zhang et al. [2024]; Gauthier [2024]; Yang et al. [2024]; Chen et al. [2024]; Ma et al. [2024]; Bouzenia et al. [2024]. While there is not a fixed definition for agent-based approaches, they generally equip LLMs with a set of tools and allow agents to iteratively and autonomously perform actions, observe feedback, and plan future steps. Example tools can include the ability to open/write/create files, search for code lines, run tests, and execute shell commands. In each attempt to solve a problem, agent-based approaches will have multiple turns, where each turn consists of performing an action. Subsequent turns depend on previous actions and the feedback information the agent receives from the environment.
@@
  • Lack of control in decision planning. In addition to using tools, current agent-based approaches also delegates the decision making process to the agents – deciding when and what action to perform. The agents decide the current action to take based on previous actions taken and the feedback provided by the environment, often with minimal checks to ensure the action taken make sense. Due to the large possible action space and feedback response, it can be extremely easy for autonomous agents to become confused and perform sub-optimal explorations. Furthermore, to solve an issue, an agent can take upwards of 30 or 40 turns which makes it extremely difficult to both understand the decisions made by the agents and also debug the exact turns where the incorrect decision is made.
  • Limited ability to self-reflect. Existing agents struggle with the capability to perform self-reflection Olausson et al. [2023]; Huang et al. [2024]. That is to say they tend to take all information/feedback and do not know how to filter out or correct irrelevant, incorrect, or misleading information Shi et al. [2023]; Zhang et al. [2023b]. For example, a common step in the agent-based approach is to reproduce an issue with a minimal test case. However, the reproduced test case may not be always correct or precise. The limited ability to self-reflect means that an incorrect step can be easily amplified and negatively affect all future decisions taken by the agent.
@@
Our work. We set out to answer this important question by building Agentless – an agentless approach to automatically solve software development problems. To solve each issue, Agentless follows a simple two phase process: localization and repair. In the localization process, Agentless employs a hierarchical process to first localize the fault to specific files, then to relevant classes or functions, and finally to fine-grained edit locations. To perform repair, Agentless takes the edit locations and generates multiple candidate patches in a simple diff format. Agentless then performs simple filtering to remove any patches that have syntax errors or cannot pass the previous tests in the repository. Finally, Agentless re-ranks all remaining patches and selects one to submit in order to fix the issue. While Agentless leverages LLMs to perform each detailed task, unlike prior complex agent-based tools, Agentless does not allow LLMs to autonomously decide future actions or operate with any complex tools. Our deliberate choice of not using agents not only allows Agentless to have a simplistic and straightforward design that can be easily understood, but also helps avoid the above mentioned limitations of LLM agents in software development. We evaluate Agentless on the popular SWE-bench Lite [swe, 2024] benchmark and demonstrate that Agentless not only achieves the highest performance (27.33%) among all open-source approaches, but it does so at a fraction of the cost!
@@
Furthermore, we performed fine-grained manual analysis on the SWE-bench Lite dataset and classified all its problems into different categories across dimensions like problem description, ground truth patch, and location information. Surprisingly, we observed that SWE-bench Lite contains problems (4.3%) with exact ground truth patch in the description, problems (9.3%) with missing critical information needed to solve the issue, and problems (4.3%) that include misleading solutions in the issue description. Recognizing these issues, we built SWE-bench Lite-$S$, which removes such problematic problems, and serves as a more rigorous benchmark to evaluate the ability to solve real-world software development problems. Overall, in an era focused on achieving top placements on leaderboards, our work highlights the overlooked potential of a simplistic, interpretable technique in autonomous software development. We hope Agentless will help reset the baseline, starting point, and horizon for autonomous software agents, and inspire future work along this crucial direction.
@@ 2 Agentless [Uncaptioned image] Approach
Figure 1 shows the overview of Agentless, consisting of two phases: localization and repair. We first take in the issue description and the existing project codebase as input. Then, we begin our hierarchical localization process by 1 turning the project codebase into a tree-like structure format that demonstrates the relative location of each file in the project. Next, 2 using this repository structure format along with the original issue description, we ask the LLM to localize and rank the top N most suspicious files that need editing to solve the issue. However, not all contents in each file need to be modified. As such, 3 we provide a skeleton for each file (i.e., a list of declaration headers of the classes and functions) and ask the LLM to output a specific list of classes and functions that we should examine more closely to fix the bug. We then provide the complete code content of the previous locations and 4 ask the LLM to finalize a smaller set of edit locations (i.e., classes, functions, or even specific lines). For the repair phase, we provide the code snippets at these edit locations together with the issue description and 5 ask the LLM to sample multiple patches to solve the issue. Next, 6 we perform a simple filtering to remove any patches with syntax errors and regression test failures, and use majority voting to rank the remaining patches. Finally, 7 Agentless selects the top-ranked patch as the final patch for submission. We now describe the steps in each of Agentless's two phases in more detail.
@@ 2.1 Localization [Uncaptioned image]
To fix or implement a new feature, the first step is to obtain the locations in the source code, as without the correct locations, it can be impossible to provide the right edits. The difficulty lies in the fact that there could be hundreds of files with thousands of lines of code each in a repository, whereas the correct locations to edit are only a few selected lines or functions. Agentless addresses this by using a simple three-step hierarchical localization process: 1) localize to selected files; 2) localize each selected files into relevant classes, functions, and variables; 3) localize to code edit locations.
@@
Localize to suspicious files. First, Agentless localizes the possible locations to specific suspicious files. Instead of providing the complete code snippet for each file in the repository, Agentless constructs a succinct representation of the repository's file and directory structure. We refer to this as the repository structure format, which begins with the root folder of the repository and organizes code files or folder names. Files and folders at the same directory level are aligned vertically, and files/folders in sub-directories are indented. We recursively traverse the entire repository to obtain the repository structure format, which will be used as input for the LLM. The repository structure format provides the necessary file paths alongside the neighboring file names to maintain organizational information in the original codebase. Agentless then inputs the processed repository structure format along with the original issue description to an LLM and requests it to identify a list of the top N suspicious files in the repository that need further inspection or modification to resolve the issue.
Figure 2: Skeleton format
@@
Localize to related elements. After obtaining the list of suspicious files to edit to solve the issue, Agentless then moves on to the second part of the localization process: localize to related elements within the suspicious files. Directly providing the complete context of all files can be large. As such, Agentless builds a compressed format of each file that contains the list of class, function, or variable declarations. We refer to this format as skeleton format, with an example shown in Figure 2. In the skeleton format, we provide only the headers of the classes and functions in the file. For classes, we further include any class fields and methods (signatures only). Additionally, we also keep comments in the class and module level to provide further information. Compared to providing the entire file context to the model, the skeleton format is a much more concise representation, especially when the file contains thousands of lines, making it impractical/costly to process all at once with existing LLMs. We provide the skeleton of all suspicious files to the LLM at one time in a single prompt, enabling the model to comprehensively analyze the pertinent information and decide the most relevant elements. Using this input, we ask the LLM to provide a list of related classes and functions that one should examine to fix the provided issue.
@@
Localize to edit locations. The previous localization step provided us with a list of related code elements. We then directly provide the code content from these elements to the model and ask it to localize to specific edit locations. Compared to using the entire file content, the input context we provide here is much smaller. With this input, we then ask the LLM to identify the final set of edit locations, specified by line numbers, functions, or classes. Our simple hierarchical localization process allows Agentless to select a set of relevant code snippets as edit locations to perform repair.
Figure 3: Search/Replace edit format

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have significantly advanced the automation of software development tasks, including code synthesis, program repair, and test generation. More recently, researchers and industry practitioners have developed various autonomous LLM agents to perform end-to-end software development tasks. These agents are equipped with the ability to use tools, run commands, observe feedback from the environment, and plan for future actions. However, the complexity of these agent-based approaches, together with the limited abilities of current LLMs, raises the following question: Do we really have to employ complex autonomous software agents? To attempt to answer this question, we build Agentless – an agentless approach to automatically resolve software development issues. Compared to the verbose and complex setup of agent-based approaches, Agentless employs a simplistic three-phase process of localization, repair, and patch validation, without letting the LLM decide future actions or operate with complex tools. Our results on the popular SWE-bench Lite benchmark show that surprisingly the simplistic Agentless is able to achieve both the highest performance (32.00%, 96 correct fixes) and low cost (\$0.70) compared with all existing open-source software agents! In fact, Agentless has already been adopted by OpenAI as the go-to approach to showcase the real-world coding performance of both GPT-4o and the new OpenAI o1 models. Furthermore, we manually classified the problems in SWE-bench Lite and found problems with exact ground truth patches or insufficient/misleading issue descriptions. As such, we construct SWE-bench Lite-$S$ by excluding such problematic issues to perform more rigorous evaluation and comparison. Our work highlights the currently overlooked potential of a simplistic, cost-effective technique in autonomous software development. We hope Agentless will help reset the baseline, starting point, and horizon for autonomous software agents, and inspire future work along this crucial direction. We have open-sourced Agentless at: https://github.com/OpenAutoCoder/Agentless
@@ 1 Introduction
Large language models (LLMs) have become the go-to default choice for code generation [Chen et al., 2021; Austin et al., 2021; Li et al., 2023; Wei et al., 2023]. State-of-the-art LLMs like GPT-4 [OpenAI, 2023] and Claude 3.5 Sonnet Anthropic [2024] have demonstrated their prowess in being able to synthesize code snippets based on given user description. However, compared to the main evaluation setting of simple, self-contained problems, applying LLMs on repository-level software engineering tasks has been understudied. Software engineering tasks like feature addition, program repair, and test generation require an in-depth understanding of not only information within files, containing thousands of lines of code, but also repository-level dependencies across files.
@@
To solve the challenging real-world software development problems from SWE-bench, inspired by the Devin AI Software Engineer dev [2024], there has been a significant body of work from both academia and industry focusing on developing agent-based approaches Zhang et al. [2024c]; Gauthier [2024]; Yang et al. [2024a]; Chen et al. [2024]; Ma et al. [2024]; Bouzenia et al. [2024]. While there is not a fixed definition for agent-based approaches, they generally equip LLMs with a set of tools and allow agents to iteratively and autonomously perform actions, observe feedback, and plan future steps Liu et al. [2024c]. Example tools can include the ability to open/write/create files, search for code lines, run tests, and execute shell commands. In each attempt to solve a problem, agent-based approaches will have multiple turns, where each turn consists of performing an action. Subsequent turns depend on previous actions and the feedback information the agent receives from the environment.
@@
  • Lack of control in decision planning. In addition to using tools, current agent-based approaches also delegate the decision-making process to the agents, allowing them to decide when and what action to perform. The agents decide the current action to take based on previous actions taken and the feedback provided by the environment, often with minimal checks to ensure the action taken make sense. Due to the large possible action space and feedback response, it can be extremely easy for autonomous agents to become confused and perform sub-optimal explorations. Furthermore, to solve an issue, an agent can take upwards of 30 or 40 turns, which makes it extremely difficult to both understand the decisions made by the agents and also debug the exact turns where an incorrect decision is made.
  • Limited ability to self-reflect. Existing agents struggle with the capability to perform self-reflection Olausson et al. [2023]; Huang et al. [2024]. That is to say they tend to take all information/feedback and do not know how to filter out or correct irrelevant, incorrect, or misleading information Shi et al. [2023]; Zhang et al. [2023]. The limited ability to self-reflect means that an incorrect step can be easily amplified and negatively affect all future decisions made by the agent.
@@
Our work. We set out to answer this important question by building Agentless – an agentless approach to automatically resolve software development issues. To solve each issue, Agentless follows a simple three phase process: localization, repair, and patch validation. In the localization process, Agentless employs a hierarchical process to first localize the fault to specific files, then to relevant classes or functions, and finally to fine-grained edit locations. Agentless's localization process make uses of both LLM-based localization as well classic information-retrieval-based localization idea Zhou et al. [2012]. To perform repair, Agentless takes the localized edit locations and generates multiple candidate patches in a simple diff format. At the same time, Agentless generates reproduction tests that can reproduce the original error and help with candidate patch selection. Finally, Agentless re-ranks all remaining patches and selects one to submit in order to fix the issue. While Agentless leverages LLMs to perform each detailed task, unlike prior complex agent-based tools, Agentless does not allow LLMs to autonomously decide future actions or operate with any complex tools. Our deliberate choice to avoid using agents not only allows Agentless to have a simplistic and straightforward design that is easy to understand, but also helps avoid the above-mentioned limitations of LLM agents in software development. We evaluate Agentless on the popular SWE-bench Lite [swe, 2024] benchmark and demonstrate that Agentless not only achieves the highest performance (32.00%) among all open-source approaches, but also does so at a fraction of the cost!
@@
Furthermore, we performed a fine-grained manual analysis on the SWE-bench Lite dataset and classified all its problems into different categories across dimensions like problem description, ground truth patch, and bug location information. Surprisingly, we observed that SWE-bench Lite contains problems with exact ground truth patch in the description (4.3%), problems with missing critical information needed to solve the issue (10.0%), and problems that include misleading solutions in the issue description (5.0%). Recognizing these issues, we built SWE-bench Lite-$S$, which removes such problematic problems, and serves as a more rigorous benchmark to evaluate the ability to solve real-world software development challenges. Overall, in an era focused on achieving top placements on leaderboards, our work highlights the overlooked potential of a simplistic, cost-effective technique in autonomous software development. We hope Agentless will help reset the baseline, starting point, and horizon for autonomous software agents, and inspire future work along this crucial direction.
Contributions. In this work, we make the following contributions:
  • An agentless approach. We propose Agentless, an agentless approach to automatically solve software development problems. Agentless leverages LLM-empowered prompting-based and embedding-based retrieval to perform hierarchical localization. During repair, Agentless samples multiple candidate patches in a simple diff format for efficient patch generation. Agentless then generates reproduction tests to verify that the issue has been resolved. Finally, Agentless leverages both regression tests and generated reproduction tests to select the final submission patch.
  • Extensive evaluation. We evaluate Agentless on the popular SWE-bench Lite dataset comparing against state-of-the-art agent-based approaches. Our results demonstrates that Agentless is able to achieve higher performance (32.00%, 96 correct fixes) than all open-source approaches, with comparably low cost as well. This shows the previously overlooked potential of a simplistic technique in autonomous software development. Additionally, Agentless has already been adopted by OpenAI as the go-to approach to showcase the real-world coding performance of GPT-4o Chowdhury et al. [2024] as well as the new o1 model family OpenAI [2024c] We further perform a rigorous ablation study to understand the effectiveness of different components of Agentless on the final performance.
  • SWE-bench Lite-$S$ benchmark. We performed manual classifications on the problems in the popular SWE-bench Lite dataset. We found that there are problematic problems with unclear, misleading issue descriptions, as well as problems that contain exact ground truth patches. To address these issues, we constructed a filtered dataset of SWE-bench Lite-$S$ that excludes such problematic issues, enabling more rigorous evaluation and comparison. This has also been confirmed recently by OpenAI, which acknowledged our benchmark and released SWE-bench Verified Chowdhury et al. [2024] along the same direction.
2 Background and Related Work
2.1 Agent-based Software Engineering
With the emergence and popularity of agent-based frameworks Xi et al. [2023], recently researchers and industry practitioners have begun developing agent-based approaches to solve software engineering tasks. Devin dev [2024] (and OpenDevin ope [2024b], open-source alternative), is one of the first end-to-end LLM agent-based framework. Devin uses agents to first perform planning based on user requirement, then allows the agent to use file editor, terminal, and web search engine tools to iteratively perform the task. SWE-agent Yang et al. [2024a] designs a custom agent-computer interface (ACI) that allows the LLM agent to interact with the repository environment with actions such as reading, editing files, and running bash commands. Aider Gauthier [2024] first provides a detailed repository map constructed with static and call graph analysis to the LLM to localize the files that require editing; then it generates a simple diff format as the editing patch and uses regression testing to verify if the patch is plausible. Moatless moa [2024] is another open-source agent tool that obtains relevant code locations by providing the agent with both code search tools as well as retrieval methods using LLM-constructed queries. Similar to Aider, Moatless also generates a simple diff format as the final submitted patch. AutoCodeRover Zhang et al. [2024c] further provides the LLM agent with specific code search APIs (e.g., searching methods in a certain class) to iteratively retrieve code context and locate the bug locations. SpecRover Ruan et al. [2024] later improves over AutoCodeRover and targets specifications (i.e., inferring the intended program behavior) by generating function summaries and also feedback messages during specific agent steps. Furthermore, SpecRover also attempts to generate reproduction tests to reproduce the original issue used to select the final patch. In addition to these highlighted examples, there has been a plethora of other agent-based approaches developed in both open-source Gauthier [2024]; rep [2024]; app [2024]; Zhang et al. [2024d] and close-source/commercial products Bouzenia et al. [2024]; Chen et al. [2024]; Ma et al. [2024]; lin [2024]; fac [2024]; ibm [2024]; ope [2024a]; Liu et al. [2024b]; ama [2024]; sup [2024]; gru [2024]; iso [2024]; men [2024]; hon [2024]; cod [2024].
Compared to these agent-based techniques, Agentless offers a simplistic, interpretable, and cost-effective solution to tackle real-world software engineering issues. Different from agent-based tools, Agentless contains well-defined stages of localization, repair, and patch validation without letting the LLM agent decide future actions or use complex tools. Agentless demonstrates for the first time that an agentless approach can achieve very competitive performance, without the additional baggage of having to provide excessive tools or model complex environment behavior/feedback.
2.2 Fault Localization and Program Repair
Fault localization (FL) Wong et al. [2016] techniques aim to identify the suspicious locations (e.g., statements or methods) in source code related to bugs. Dynamic FL techniques mainly include spectrum-based FL (SBFL) Jones and Harrold [2005]; Abreu et al. [2007, 2009] and mutation-based FL (MBFL) Papadakis and Le Traon [2015]; Moon et al. [2014]; Lou et al. [2020]. SBFL typically computes source code locations primarily covered by failing tests as more suspicious than locations primarily covered by passing tests. MBFL further improves upon that to additionally consider the impact of each source code location on the test outcomes (measured using mutation testing Papadakis et al. [2019]). Besides dynamic techniques, researchers have also proposed to directly leverage information retrieval (IR) techniques Singhal et al. [2001] for static FL. Such IR-based techniques Wang et al. [2015]; Saha et al. [2013]; Wang and Lo [2014] formulate FL as a search problem and compare the textual similarity between code elements and the bug report (i.e., query). Moreover, learning-based techniques have also been proposed to leverage machine learning to combine multiple sources of dynamic/static information, including DeepFL Li et al. [2019], FLUCCS Sohn and Yoo [2017], and TRANSFER Meng et al. [2022]. Recently, researchers have proposed LLM-based FL Yang et al. [2024b]; Wu et al. [2023]; Qin et al. [2024]; Kang et al. [2024], which leverages the powerful code and natural language understanding of modern LLMs to directly localize bugs. Meanwhile, most such LLM-based techniques either cannot perform repository-level FL due to the limited context window of LLMs Yang et al. [2024b], or rely on complicated/costly agentic design to navigate through the codebase Kang et al. [2024]; Qin et al. [2024]. In contrast, Agentless employs a simplistic hierarchical FL process (based on both LLMs and IR) to efficiently compute the fine-grained edit locations.
After localizing the bug, the next step is to perform repair. Automated program repair Gazzola et al. [2019] (APR) has been widely studied to automatically generate patches for bugs. Traditional APR techniques can be categorized as template-based Liu et al. [2019]; Ghanbari et al. [2019], heuristic-based Le Goues et al. [2012]; Le et al. [2016], and constraint-based Mechtaev et al. [2016]; Long and Rinard [2015] tools. While effective, traditional APR tools suffer from scalability issues and are limited by their patch variety. As such, researchers have proposed learning-based APR tools either by training NMT (neural machine translation) models Jiang et al. [2021]; Li et al. [2020]; Chen et al. [2019]; Zhu et al. [2021] or using pre-trained LLMs to perform repair Xia and Zhang [2022]; Xia et al. [2023a]; Kolak et al. [2022]; Zhang et al. [2024a]. Specifically, LLM-based APR tools, which sample multiple candidate patches per bug, have been shown to be the state-of-the-art due to the powerful coding capability of modern LLMs Xia et al. [2023a]. More recently, agent-based APR techniques have also been proposed Xia and Zhang [2023]; Bouzenia et al. [2024]; Hidvégi et al. [2024]; Chen [2024]; Zhang et al. [2024b]. Inspired by existing LLM-based APR tools, Agentless samples multiple candidate patches per issue to maximize the chance of generating a correct fix. Different from most LLM-based APR techniques, Agentless generates patches using a simple diff format Gauthier [2024] to avoid generating the complete code and instead focus on producing cost-efficient small edits, increasing the reliability and accuracy of patch generation (less chances for hallucination). Furthermore, different from the simplistic bugs studied in most prior work Xia et al. [2023a], Agentless targets complex repository-level issues spanning multiple locations.
2.3 LLM-based Test Generation
In addition to localizing and repairing bugs, another research area that has been adopting LLM is test generation. One area of test generation is fuzz testing Zeller et al. [2019], also known as fuzzing, to generate large amounts of inputs in order to expose bugs in systems. Researchers have applied LLMs to perform fuzzing in domains such as DL libraries Deng et al. [2023, 2024], OS Kernel Yang et al. [2023b]; Oliinyk et al. [2024], compilers Xia et al. [2024]; Yang et al. [2023a]; Ou et al. [2024], network protocols Meng et al. [2024], and mobile applications Liu et al. [2024a]. LLM-based fuzzers have demonstrated their impact by detecting many bugs not found by traditional fuzzers as well as unlocking new fuzzing domains. Besides fuzzing, researchers have also proposed to leverage LLMs for unit test generation to test individual software units (e.g., methods/classes) Liu et al. [2024c], such as CodeMosa Lemieux et al. [2023], ChatTester Yuan et al. [2024], TestPilot Schäfer et al. [2023], and CoverUp Pizzorno and Berger [2024].
Bug reproduction is a critical step in investigating bug reports Jin and Orso [2012], and is integrated into many recent software engineering agents ope [2024b]; Ruan et al. [2024]; hon [2024]; men [2024]; fac [2024]; Yang et al. [2024a]; Arora et al. [2024]. For example, SpecRover Ruan et al. [2024] begins by generating a test to reproduce the issue described in the bug report; the test then guides the context retrieval and patching process.Unlike such agent-based approaches, which generate a reproduction test and rely on an LLM agent to decide whether the test is correct, Agentless simply executes multiple sampled tests and verifies if the execution results indicate the issue has been reproduced.
@@ 3 Agentless [Uncaptioned image] Approach
Figure 1 shows the overview of Agentless, consisting of three phases: localization, repair, and patch validation. We first take in the issue description and the existing project codebase as input. Then, we begin our hierarchical localization process by turning the project codebase into a tree-like structure that illustrates the relative location of each file in the project 1 . Next, using this repository structure along with the original issue description, we prompt the LLM to localize and rank the top N most suspicious files that likely require editing to solve the issue 2 . Since our repository structure format does not contain detailed source code information, we additionally retrieve files with most relevant code snippets with the issue description using embedding-based retrieval 3 . We then combine the retrieved files with the LLM-localized files to obtain the final list of suspicious files. However, not all contents in each file need to be modified. As such, we provide a skeleton for each file (i.e., a list of declaration headers of the classes and functions) and ask the LLM to output a specific list of classes and functions that we should examine more closely to fix the bug 4 . We then provide the complete code content of the previous locations and ask the LLM to finalize a smaller set of edit locations (i.e., classes, functions, or even specific lines) 5 . For the repair phase, we provide the code snippets at these edit locations together with the issue description and prompt the LLM to sample multiple patches to solve the issue 6 . Next, we enter the patch validation phase, where we first ask the LLM to sample multiple reproduction tests that aim to replicate the original issue 7 , and then select the optimal one based on actual execution results on the original codebase 8 . Agentless uses the reproduction test along with existing regression tests for patch ranking/selection 9 . Finally, Agentless selects the top-ranked patch as the final patch for submission 10 . We now describe the steps in each of Agentless's phases in more detail.
@@ 3.1 Localization [Uncaptioned image]
To fix or implement a new feature, the first step is to obtain the locations in the source code, as without the correct locations, it can be impossible to provide the right edits. The difficulty lies in the fact that there could be hundreds of files with thousands of lines of code each in a repository, whereas the correct locations to edit are only a few selected lines or functions. Agentless addresses this by using a simple three-step hierarchical localization process: 1) localize to suspicious files; 2) localize each selected files into relevant classes, functions, and variables; 3) localize to code edit locations.
3.1.1 Localize to suspicious files.
@@
First, Agentless narrows down potential locations to specific suspicious files. Instead of providing the complete code snippet for each file, Agentless constructs a concise representation of the repository's file and directory structure, similar to the Linux tree command. We refer to this as the repository structure format, which begins with the root folder of the repository and organizes code files or folder names. Files and folders at the same directory level are aligned vertically, and files/folders in sub-directories are indented. We recursively traverse the entire repository to obtain the structure, which will be used as input for the LLM. The repository structure format provides the necessary file paths alongside the neighboring file names to maintain organizational information in the original codebase. Agentless then inputs the processed repository structure along with the original issue description to an LLM, and requests it to identify a list of the top N suspicious files that need further inspection or modification to resolve the issue.
To compliment the prompting-based localization (using file names only), Agentless also uses a simple embedding-based retrieval method to identify additional suspicious files. However, instead of embedding all files in the repository, Agentless first filters out irrelevant folders. This is done by providing the previously described repository structure and asking the LLM to produce a list of irrelevant folders that do not need to be further inspected or modified to resolve the issue. After removing all files from these irrelevant folders, Agentless divides each remaining file into chunks of code segments and computes the embedding for each chunk using an embedding model. Agentless then embeds the original issue description (i.e., the query) and computes the cosine similarity between the resulting query embedding and each chunk embedding to retrieve a list of relevant files that contain code segments with the highest similarity to the query. Finally, Agentless combines the files obtained via prompting with those retrieved via embedding by selecting top N most common files localized by both, resulting in a final list of relevant files.
Figure 2: File skeleton format.
3.1.2 Localize to related elements.
@@
After obtaining the list of suspicious files, Agentless proceeds to the second part of the localization process: localize the related elements within these files. Directly providing the complete context of all files can be large. As such, Agentless builds a compressed format of each file that contains the list of class, function, or variable declarations. We refer to this format as the skeleton format, with an example shown in Figure 2. In the skeleton format, we provide only the headers of the classes and functions in the file. For classes, we further include any class fields and methods (signatures only). Additionally, we also keep comments in the class and module level to provide further information. Compared to providing the entire file context to the model, the skeleton format is a much more concise representation, especially when the file contains thousands of lines, making it impractical/costly to process all at once with existing LLMs. We provide the skeleton of all suspicious files to the LLM at one time in a single prompt, enabling the model to comprehensively analyze the pertinent information and decide the most relevant elements. Using this input, we prompt the LLM to provide a list of related classes and functions that one should examine to fix the provided issue.
Figure 3: Search/Replace edit format.
3.1.3 Localize to edit locations.
@@
The previous localization step provided us with a list of related code elements; since we localize top N suspicious files, these localized related code elements could be from different files. We now directly provide the code content from these elements to the LLM and ask it to localize specific edit locations. Compared to using the entire file, the input context here is much smaller. With this input, we then ask the LLM to identify the final set of edit locations, specified by line numbers, functions, or classes. Our simple hierarchical localization allows Agentless to select a set of relevant code snippets as edit locations for repair.